Identify the line of symmetry, the vertex, the y - intercept, and the x - intercept
f(x) = x^2 - 10x + 9
PLEASE HELP !!!

Answer by Cromlix(4381)

f(x) = x^2 - 10x + 9
Factorised:
(x - 1)(x - 9) = 0
Cuts x axis at (1,0) and (9,0)
Cuts y axis when x = 0
x^2 - 10x + 9
(0)^2 - 10(0) + 9
y = 9
Cuts y axis at (0, 9)
Line of symmetry is half way between
x = 1 and x = 9
8 units/ 2
= 4 units
at x = 5.
Vertex = Put x = 5 into f(x) to get y coord.
y = 5^2 - 10*5 + 9
y = 25 - 50 + 9
y = -16
So vertex at (5, -16)
Hope this helps.
;-)